A truck with 0.375 m radius tires travels at 35.5 m/s. At how many radians per second are the tires rotating?
___ rad/s

What is this value in rpm?
____ rpm
Answer #1

Solution)

Part a)

Given, Radius, r=0.375 m

Speed, v= 35.5 m/s

Now, angular velocity, w=v/r= 35.5/0.375= 94.66 rad/s (Ans)

===========

In RPM

=94.66*(1/2pi)*60= 904.39 =904 rpm (Ans)

================